# Mediation analysis in longitudinal intervention studies with an ordinal treatment-dependent confounder

### v1

@@ 3.3 Identification of the causal effects
The first three assumptions translate to no unmeasured confounding between $A$, $U$, $M$ and $T$. The assumptions 4 and 5 are cross-world independencies, essentially requiring that there exist no further confounders between $U$, $M$ and $T$ which are affected by the treatment. Under these assumptions, the direct and indirect effects (1) can be identified up to a term containing the joint distribution of $(U_{a},U_{a^{*}})$. We will show below that in the case of a trichotomous treatment-dependent confounder $U$ and assuming the monotonicity of the treatment effect on $U$ will further allow the identification of the estimands up to a scalar sensitivity parameter.
Let $m^{\dagger}$ and $u^{\dagger}$ denote some reference trajectory for the mediator and a reference value for $U$, respectively. Following Tchetgen Tchetgen and VanderWeele (2014), the expected outcome, given $A$, $M$, $U$, $\boldsymbol{W}$ and $R_{0}$, can be decomposed into the main effects of $M$ and $U$, the interaction between $M$ and $U$ and a reference level:
@@
Tchetgen Tchetgen and VanderWeele showed that assuming no additive interaction of $M$ and $U$ on $T$, or, in the case of a binary $U$, monotonicity of the effect of $A$ on $U$, would be enough to identify $\delta$, and therefore also the direct and indirect effects, from empirical data (Tchetgen Tchetgen and VanderWeele, 2014). The ‘no additive interaction’ assumption would imply $\Delta_{DE}=\Delta_{IE}=\delta=0$, and subsequently $DE=DE^{(r)}$ and $IE=IE^{(r)}$. The monotonicity assumption of the effect of $A$ on $U$ means that an individual cannot have a worse value of $U$ under treatment than they would have had under no treatment. If $U$ is binary, the joint probability of $U_{a}$ and $U_{a^{*}}$ becomes fully determined by the marginal probabilities, which can be estimated from the observed data.
We now extend the monotonicity assumption to a trichotomous $U$ and show that the joint probability, conditionally on the marginals, is identified up to a scalar sensitivity parameter restricted by the marginals. Supposing that $U$ is trichotomous, taking values $U\in\{0,1,2\}$, and the treatment effect on $U$ is monotonic. The stratum-specific joint probability of $(U_{a},U_{a^{*}})$ can then be represented in terms of six probability parameters:
@@ 0 $p_{00}$ $p_{01}$ $p_{02}$ $\Phi_{4}$
The admissible values for $p_{11}$ can now be expressed in terms of a sensitivity parameter $\rho=(p_{11}-p_{min})/(p_{max}-p_{min})\in[0,1]$, describing where $p_{11}$ lies within its supported interval. For a fixed $\rho$, the term $\delta$ is identified and its range of possible values can be calculated by varying the sensitivity parameter $\rho$ within the $[0,1]$ interval. Furthermore, the consistency of the monotonicity assumption with empirical data can be assessed by estimating the marginal probabilities and verifying whether the condition $1-\Phi_{3}-\Phi_{4}\leq\min\{\Phi_{2},\Phi_{5}\}$ holds.
@@ 3.4 Parametric models
Under the identifying assumptions of Section 3.3, the direct and indirect effects can be expressed in terms of the observed data. Although the effects and their corresponding empirical expressions are derived nonparametrically, the components forming the empirical expressions are estimated parametrically. In this section we describe the parametric models we used to estimate the terms in the expressions (2) for the direct and indirect effects.
@@
where $\boldsymbol{X}_{i}=(A_{i},I(U_{i}=1),I(U_{i}=2),A_{i}I(U_{i}=1),A_{i}I(U_{i}=2))^{\prime}$ and $\boldsymbol{W}_{i}$ is a vector containing the baseline covariates (age, sex, smoking status and lifestyle score). The terms $B_{k}$ are population-level basis functions for natural cubic splines with outer knots placed at the baseline and 10 years and three inner knots placed at 1, 3 and 5 years since study onset. Similarly, $B^{r}_{k}$ are basis functions corresponding to individual-level random effects, with two inner knots placed at years 1 and 5. The $\epsilon_{ij}$ are normally distributed mutually independent error terms with variance $\sigma^{2}$ and the random effects $(R_{i0},R_{i1},R_{i2},R_{i3})^{\prime}$ have a multinormal distribution with mean zero and a full covariance matrix $\Sigma$. Note that the model includes non-time-dependent effects of the treatment and treatment-dependent confounder on the outcome. This entails the assumptions that the treatment-dependent confounder is affected by the treatment without delay, and it in turn affects the mediator trajectory without delay.
@@
The choice of the functional form of $g(\cdot)$ depends on the biological mechanism by which the mediator is assumed to affect the outcome. Here, we choose a three-year legacy parameterisation, i.e. $g(M_{i}(\cdot),t)=\int_{0}^{3}\big{[}M_{i}(v)-M_{i}(0)\big{]}\,\text{d}v$ for all $t\geq 0$, implying that the hazard at any time is affected by the cumulative change of the latent trajectory over the first three years since the baseline. We return to discuss the interpretation of this choice in the Discussion. For comparison, we also consider a current change parameterisation where the hazard at time $t$ is affected by the change by time $t$ of the level of the latent trajectory since the baseline, i.e. $g(M_{i}(\cdot),t)=M_{i}(t)-M_{i}(0)$.
@@ 3.5 Estimation
We employed a joint modelling framework to estimate the parametric models of the mediator trajectory and the time-to-event outcome. Since these models use only the observed values of the treatment-dependent confounder, the model for the treatment-dependent confounder was estimated separately. A joint model for the longitudinal and time-to-event outcomes comprises specifying submodels for both outcomes and linking them via some association structure, thus allowing incorporating any information shared between the two outcomes (Gould et al., 2015; Ibrahim et al., 2010; Hickey et al., 2018; Papageorgiou et al., 2019). The association structure was here induced by including a function of the longitudinal mediator trajectory into the linear predictor of the survival submodel and also assuming a random effect which is shared between the mediator trajectory and the time-to-event outcome enabling adjustment for a latent confounder.
@@ 4 Application
The DPS followed a cohort of 522 individuals, of which 265 were randomised to the intervention group and 257 to the control group. After removing individuals with missing values on the baseline covariates ($n=19$), the final sample consisted of 503 individuals with 254 people in the intervention group and 249 in the control group. The median number of clinical visits per person was 11 (interquartile range 6–13) in the intervention group and 9 (IQR 4–13) in the control group. The intervention group contributed a total of 3505 person years during which 166 T2D cases were observed with a restricted mean survival time of 11.2 years, while the control group contributed a total of 2865 person years and 168 T2D cases with a restricted mean survival time of 9.5 years.
The study endpoint was the diagnosis of T2D, ascertained either at any of the study visits or inferred through the register data. As the clinical study visits were arguably the more reliable source, we used the first diagnosis made at the study visits as the primary endpoint and considered the register data for each individual only after their last study visit. Figure 2($a$) shows the cumulative incidences of T2D from both sources. The clinical risk set refers to the number of individuals in the risk set having not yet made their last study visit, whereas the register risk set is the number of individuals in the risk set being followed through the drug registers. The cumulative incidence curves based on the study visits and the drug registers grow reasonably closely in proportion to the number of individuals, implying that any bias due to uneven sensitivity of T2D detection was unlikely.
For each study participant, the follow-up started at the baseline visit and terminated at the event of T2D diagnosis, death or end of follow-up at the end of 2018. The dates and causes of deaths were available through the Finnish Cause of Death Register. Figure 2 (b) shows the cumulative risks for the competing events of T2D diagnosis, death with a possibly T2D-related cause, namely cardio- and cerebrovascular complications, and other causes. As the proportion of potentially T2D-related deaths appears negligible, treating all deaths as uninformative right censorings was deemed justified.
@@ 4.1 Model fit
Figure 3(a) shows the observed and estimated mean BMI trajectories in the two treatment groups. Figures 3(b) to 3(d) show the Kaplan-Meier curves and estimated survival functions, illustrating separately the effects of the treatment, change in BMI over the first three years and lifestyle score on avoiding T2D. In these plots, the cumulative change in BMI over the first three years was categorised into tertiles of the individual level point estimates obtained from the model. The survival functions were then computed as the means of the estimated survival functions for each of the three groups. The plots are based on the model using the three-year legacy parameterisation
The mean BMI trajectory in the intervention group (Figure 3(a)) exhibits a decrease in the early years since the start of the intervention, after which the mean trajectory gradually rebounds close to the trajectory of the control group. The estimated curves reproduce the observed average trajectories quite well, although the shape is slightly different in the intervention group over the first two years. The Kaplan-Meier curves indicate lower T2D risk for those in the intervention group (Figure 3(b)), in the higher lifestyle score groups (Figure 3(d)) or having decreased their BMI more over the first three years (Figure 3(c)). The estimated survival functions agree closely with the Kaplan-Meier curves. However, with the early BMI change the differences in the estimated survival functions are less pronounced than is evident from the Kaplan-Meier curves, which might indicate some lack of fit in the model with respect to the relationship between the early BMI change and T2D hazard.
@@ 4.2 Total, direct and indirect effects
0.75 $0.598$ $(-1.72,2.24)$ $0.972$ $(0.138,2.10)$ $1.57$ $(-0.082,2.95)$ $1.31$ $(0.059,2.81)$ $0.304$ $(-0.083,0.746)$ $1.61$ $(0.313,3.09)$
1 $0.595$ $(-1.72,2.24)$ $0.974$ $(0.137,2.10)$ $1.57$ $(-0.082,2.95)$ $1.31$ $(0.060,2.81)$ $0.302$ $(-0.091,0.747)$ $1.61$ $(0.313,3.09)$
Table 1 shows the posterior means and $95\%$ credible intervals (CI) of the direct, indirect and total effects with different choices of the sensitivity parameter $\rho$. The estimated effects remained practically unaffected by the sensitivity parameter. We therefore discuss the results with the choice $\rho=0.5$.
@@
As it could be considered a plausible a priori assumption, that the effects of the intervention cannot be negative, the results are also shown by discarding posterior draws ($23\%$ of the total MCMC samples under the three-year legacy and $8\%$ under the current change parameterisation) giving a negative estimate for the direct or the indirect effect (Table 2). Under such a restriction, the proportion mediated, i.e. $IE/TE$, becomes a well defined quantity and describes the proportion of the total effect that may be attributed to the indirect mechanism. With the three-year legacy parameterisation, the direct and the total effects were larger, while the indirect effect was slightly smaller. The estimated proportion mediated was roughly half, however, the $95\%$ credible interval extended over the entire $[0,1]$ interval. With the current change parameterisation there was little change in results compared with the unrestricted analysis as only $8\%$ of the posterior draws were discarded. The estimated proportion mediated was $20\%$ with a $95\%$ credible interval from $2\%$ to $64\%$.
Table 2: The posterior means and $95\%$ posterior intervals (in parentheses) of the direct ($DE$), indirect ($IE$) and total ($TE$) effects under the two alternative functionals after discarding Markov chain Monte Carlo samples giving negative treatment effect estimates. The sensitivity parameter was set at $\rho=0.5$. Results after discarding posterior draws resulting in negative effect estimates. The sensitivity parameter was set at $\rho=0.5$.
@@
4.3 Assessment of the monotonicity assumption
Figure 4: $(a)$ Proportions of Markov chain Monte Carlo samples failing the monotonicity assumption in the 54 strata defined by the covariates. The largest proportions of failures occurred in strata with low numbers of individuals. (b) The average posterior boundaries within which the free parameter in the joint distribution of the two counterfactual treatment-dependent confounders ($p_{11}$) was constrained by the monotonicity assumption in each stratum.
The agreement of the monotonicity assumption with the data was assessed by computing the estimated marginal probabilities for the two counterfactual treatment-dependent confounders ($U_{a}$ and $U_{a^{*}}$) for each strata formed by the baseline covariates, and checking whether the marginals satisfied the condition required by the monotonicity assumption. Out of the 8000 MCMC samples, only 205 ($2.6\%$) contained one or more strata for which the marginals were not consistent with the monotonicity assumption. Figure 4$(a)$ shows the proportion of these failures for each stratum and the number of individuals belonging to each in the data. The strata with the highest proportion of failures were ones with low coverage in the data, which seems reasonable since those would be expected to be the least trustworthy.

### v2

@@ 3.3 Identification of the causal effects
The first three assumptions translate to no unmeasured confounding between $A$, $U$, $M$ and $T$. The assumptions 4 and 5 are cross-world independencies, essentially requiring that there exist no further confounders between $U$, $M$ and $T$ which are affected by the treatment. In addition, we take the missing data mechanism to be missing at random, i.e. the right censorings of the survival process and drop-outs in the longitudinal follow-ups are considered non-informative. Under these assumptions, the direct and indirect effects (1) can be identified up to a term containing the joint distribution of $(U_{a},U_{a^{*}})$. We will show below that in the case of a trichotomous treatment-dependent confounder $U$ and assuming the monotonicity of the treatment effect on $U$ will further allow the identification of the estimands up to a stratum-specific sensitivity parameter.
Let $m^{\dagger}$ and $u^{\dagger}$ denote some reference trajectory of the mediator and a reference value of $U$, respectively. Following Tchetgen Tchetgen and VanderWeele (2014), the expected outcome, given $A$, $M$, $U$, $\boldsymbol{W}$ and $R_{0}$, can be decomposed into four terms representing the main effects of $M$ and $U$, the interaction between $M$ and $U$, and a reference level:
@@
Tchetgen Tchetgen and VanderWeele showed that assuming no additive interaction of $M$ and $U$ on $T$, or, in the case of a binary $U$, monotonicity of the effect of $A$ on $U$, would be enough to identify $\delta$ and, therefore, also the direct and indirect effects from empirical data (Tchetgen Tchetgen and VanderWeele, 2014). The ‘no additive interaction’ assumption would imply $\Delta_{DE}=\Delta_{IE}=\delta=0$, and subsequently $DE=DE^{(r)}$ and $IE=IE^{(r)}$. The monotonicity assumption of the effect of $A$ on $U$ means that an individual cannot have a worse value of $U$ under treatment than they would have had under no treatment. If $U$ is binary, the joint probability of $U_{a}$ and $U_{a^{*}}$ becomes fully determined by the marginal probabilities, which can be estimated from the observed data.
We now extend the monotonicity assumption to a trichotomous $U$ and show that, in each of the strata defined by the baseline covariates, the joint probability is identified up to a sensitivity parameter. For convenience, we omit denoting the stratum in what follows. Suppose that $U$ is trichotomous, taking values $U\in\{0,1,2\}$, and the treatment effect on $U$ is monotonic. The stratum-specific joint probability of $(U_{a},U_{a^{*}})$ can then be represented in terms of six probability parameters:
@@ 0 $p_{00}$ $p_{01}$ $p_{02}$ $\Phi_{4}$
The admissible values for $p_{11}$ can now be expressed in terms of a sensitivity parameter $\rho=(p_{11}-p_{min})/(p_{max}-p_{min})\in[0,1]$, describing where $p_{11}$ lies within its supported interval. For a fixed $\rho$, the term $\delta$ is identified and its range of possible values can be calculated by varying the sensitivity parameter $\rho$ within the $[0,1]$ interval. The lower and upper bounds for the direct and indirect effects can be found by minimising and maximising $\delta$ within each stratum. Since each $\delta$ is linear in the probabilities $p_{ij}$, its minimum and maximum values for a given stratum are found at the opposite ends of the $\rho$ interval. Furthermore, (3) implies that $1-\Phi_{3}-\Phi_{4}\leq\min\{\Phi_{2},\Phi_{5}\}$ is a necessary (but not sufficient) condition for the monotonicity assumption to hold. The marginal probabilities $\Phi$ can be estimated from data and thus allow one to assess based on empirical data whether the monotonicity assumption can be considered viable. Generalisation to the case with a treatment-dependent confounder with arbitrary number of levels is given in the online supplementary material.
@@ 3.4 Parametric models
Under the identifying assumptions of Section 3.3, the direct and indirect effects can be expressed in terms of the observed data. Although the effects and their corresponding empirical expressions were derived nonparametrically, the components forming the empirical expressions were estimated parametrically. In this section we describe the parametric models we used to estimate the terms in the expressions (2) for the direct and indirect effects.
@@
where $\boldsymbol{X}_{i}=(A_{i},I(U_{i}=1),I(U_{i}=2),A_{i}I(U_{i}=1),A_{i}I(U_{i}=2))^{\prime}$ and $\boldsymbol{W}_{i}$ is a vector containing the baseline covariates (age, sex, smoking status and lifestyle score). The terms $B_{k}$ are population-level basis functions for natural cubic splines with outer knots placed at the baseline and 10 years and three inner knots placed at 1, 3 and 5 years since study onset. Similarly, $B^{r}_{k}$ are basis functions corresponding to individual-level random effects, with two inner knots placed at years 1 and 5. The $\epsilon_{ij}$ are normally distributed mutually independent error terms with variance $\sigma^{2}$ and the random effects $(R_{i0},R_{i1},R_{i2},R_{i3})^{\prime}$ have a multinormal distribution with mean zero and a full covariance matrix $\Sigma$. Note that the model includes time-independent effects of the treatment and treatment-dependent confounder on the outcome. This entails the assumptions that the treatment-dependent confounder is affected by the treatment without delay, and it in turn affects the mediator trajectory without delay.
@@
The choice of the functional form of $g(\cdot)$ should depend on the biological mechanism by which the mediator is assumed to affect the outcome. Here, we choose a three-year legacy parameterisation, i.e. $g(M_{i}(\cdot),t)=\int_{0}^{3}\big{[}M_{i}(v)-M_{i}(0)\big{]}\,\text{d}v$ for all $t\geq 0$, implying that the hazard at any time is affected by the cumulative change of the latent trajectory over the first three years since the baseline. We return to discuss the interpretation of this choice in the Discussion. For comparison, we also consider a current change parameterisation where the hazard at time $t$ is affected by the change by time $t$ of the level of the latent trajectory since the baseline, i.e. $g(M_{i}(\cdot),t)=M_{i}(t)-M_{i}(0)$.
@@
The above model includes only the main effects of each predictor and could thus be considered relatively inflexible. As the research question is concerned with the mediating mechanism of the treatment, we also considered models with interaction terms between the treatment and the baseline covariates to allow the treatment effect on the treatment-dependent confounder to differ among the baseline covariate strata. Model comparison was carried out to determine whether any of the more flexible models should be favoured against model (7).
@@ 3.5 Estimation
We employed a joint modelling framework to estimate the parametric models of the mediator trajectory and the time-to-event outcome. Since these models rely on only the observed values of the treatment-dependent confounder, the model for the treatment-dependent confounder was estimated separately. Pareto smoothed importance sampling leave-one-out cross-validation (Vehtari et al., 2015) was used to compare the treatment-dependent confounder model against its more flexible variations.
A joint model for the longitudinal and time-to-event outcomes comprises specifying submodels for both outcomes and linking them via some association structure, thus allowing incorporating any information shared between the two outcomes (Gould et al., 2015; Ibrahim et al., 2010; Hickey et al., 2018; Papageorgiou et al., 2019). The association structure was here induced by including a function of the longitudinal mediator trajectory into the linear predictor of the survival submodel and also assuming a random effect which is shared between the mediator trajectory and the time-to-event outcome enabling adjustment for a latent confounder.
@@ 4 Application
The DPS originally followed a cohort of 522 individuals, with 265 randomised to the intervention group and 257 to the control group. After excluding individuals with missing values on the baseline covariates ($n=19$), the final sample consisted of 503 individuals with 254 people in the intervention group and 249 in the control group. The median number of clinical visits per person was 11 (interquartile range 6–13) in the intervention group and 9 (IQR 4–13) in the control group. The intervention group contributed a total of 3505 person years during which 166 T2D cases were observed with a restricted mean survival time of 11.2 years. The control group contributed a total of 2865 person years and 168 T2D cases with a restricted mean survival time of 9.5 years.
The study endpoint was the diagnosis of T2D, either ascertained at any of the study visits or inferred from the register data. Since the clinical study visits were considered the more reliable source, we used the first diagnosis made at the study visits as the primary endpoint and considered the register data only after their last study visit for each individual. Figure 2($a$) shows the cumulative incidences of T2D from both sources. The clinical risk set refers to the number of individuals in the risk set having not yet made their last study visit, whereas the register risk set is the number of individuals in the risk set being followed through the drug registers. The cumulative incidence curves based on the study visits and the drug registers grow reasonably closely in proportion to the number of individuals, implying that any bias due to uneven sensitivity of T2D detection was unlikely.
For each study participant, the follow-up started at the baseline visit and terminated at the event of T2D diagnosis, death or end of follow-up at the end of 2018. The dates and causes of deaths were obtained from the Finnish Cause of Death Register. Figure 2 (b) shows the cumulative risks for the competing events of T2D diagnosis, death with a potentially T2D-related cause, i.e. cardiovascular and cerebrovascular complications, and death from other causes. As the proportion of potentially T2D-related deaths appears negligible, treating all deaths as uninformative right censorings was deemed justified. With the availability of the register data after the clinical follow-up, deaths were the only source of censoring in the data.
@@ 4.1 Model fit
Figure 3(a) shows the observed and estimated mean BMI trajectories for the two treatment groups. Figures 3(b) to 3(d) display the Kaplan-Meier curves and estimated survival functions, separately illustrating the effects of the treatment, change in BMI over the first three years and lifestyle score on avoiding T2D. In these plots, the cumulative change in BMI over the first three years was categorised into tertiles of the individual level point estimates obtained from the model. The survival functions were then computed as the means of the estimated survival functions for each of the three groups. These plots are based on the model using the three-year legacy parameterisation
The mean BMI trajectory in the intervention group (Figure 3(a)) exhibits a decrease in the early years since the start of the intervention, after which the mean trajectory gradually rebounds close to the trajectory of the control group. The estimated curves replicate the observed average trajectories reasonably well, although the shape slightly differs in the intervention group over the first two years. The Kaplan-Meier curves indicate a lower T2D risk for individuals in the intervention group (Figure 3(b)), in the higher lifestyle score groups (Figure 3(d)) or having decreased their BMI more over the first three years (Figure 3(c)). The estimated survival functions agree closely with the Kaplan-Meier curves. However, with the early BMI change the differences in the estimated survival functions are less pronounced than is evident from the Kaplan-Meier curves, which might indicate some lack of fit in the model with respect to the relationship between the early BMI change and T2D hazard.
@@
4.2 Model comparison and assessment of the monotonicity assumption
In addition to model (7) of the treatment-dependent confounder, we considered more flexible models including interactions of the treatment with each baseline covariate all at the same time or each one separately. The model comparison showed no discernible difference between the model’s performances and so we chose the most parsimonious one, i.e., model (7) (see online supplement).
Figure 4: $(a)$ Proportions of Markov chain Monte Carlo samples failing the monotonicity condition in the 54 strata defined by the covariates under the selected model. The largest proportions of failures occurred in strata with low numbers of individuals. (b) The average posterior boundaries within which the free parameter in the joint distribution of the two counterfactual treatment-dependent confounders ($p_{11}$) was constrained by the monotonicity assumption in each stratum.
The monotonicity condition under the selected model was assessed from the empirical data by computing the estimated marginal probabilities for the two counterfactual treatment-dependent confounders ($U_{a}$ and $U_{a^{*}}$) for each of the 54 strata formed by the baseline covariates and checking whether the marginals satisfied the necessary condition for the monotonicity assumption. Out of the 8000 MCMC samples, only 205 ($2.6\%$) contained one or more strata for which the marginals were not consistent with the monotonicity assumption. Figure 4$(a)$ shows the proportion of these failures for each stratum and the number of individuals belonging to each in the data. The strata with the highest proportion of failures were the ones with low coverage in the data, which seems reasonable since those would be expected to be the least trustworthy. The results of the model comparison and investigating the monotonicity assumption under the alternative models can be found at the online supplementary material.
@@ 4.3 Total, direct and indirect effects
max $0.689$ $(-1.57,2.34)$ $1.05$ $(0.217,2.13)$ $1.57$ $(-0.082,2.95)$ $1.37$ $(0.116,2.87)$ $0.369$ $(-0.018,0.814)$ $1.61$ $(0.313,3.09)$
Table 1 shows the posterior means and $95\%$ credible intervals (CI) of the direct, indirect and total effects by choosing the stratum specific sensitivity parameters so that the mediational effects were either minimised or maximised. In addition, the value $\rho=0.5$ was used homogeneously in all strata. The estimated effects remained virtually the same regardless of the choice. We therefore discuss the results with the choice $\rho=0.5$.
@@
As it could be considered a plausible a priori assumption, that the effects of the intervention cannot be negative, the results are also shown by discarding posterior draws ($23\%$ of the total MCMC samples under the three-year legacy and $8\%$ under the current change parameterisation) giving a negative estimate for the direct or the indirect effect (Table 2). Under such a restriction, the proportion mediated, i.e. $IE/TE$, becomes a well defined quantity and describes the proportion of the total effect that may be attributed to the indirect mechanism. With the three-year legacy parameterisation, the direct and the total effects were larger, while the indirect effect was slightly smaller. The estimated proportion mediated was roughly half, however, and the $95\%$ credible interval extended over the entire $[0,1]$ interval. With the current change parameterisation there was little change in results compared with the unrestricted analysis as only $8\%$ of the posterior draws were discarded. The estimated proportion mediated was $20\%$ with a $95\%$ credible interval from $2\%$ to $64\%$.
Table 2: The posterior means and $95\%$ posterior intervals (in parentheses) of the direct ($DE$), indirect ($IE$) and total ($TE$) effects under the two alternative functionals after discarding Markov chain Monte Carlo samples giving negative treatment effect estimates. The sensitivity parameter was set at $\rho=0.5$.
@@
4.4 Sensitivity analyses
For sensitivity analyses, we used the approach of Miles et al. (2017) to find lower and upper bounds for the estimates relaxing the monotonicity assumption, i.e., optimising the expressions with respect to a joint probability matrix $(U_{a^{*}},U_{a})$ without the constraints implied by the monotonicity assumption. We considered three scenarios: using either (a) the full data and the same model for $U$ as in the primary analyses; (b) the same model for $U$ but removing current smokers as a potential outlier group from the data; or (c) the full data and the model for $U$ with all treatment-covariate interactions included. The lower and upper bounds for the estimates of the direct and indirect effects and their $95\%$ credible intervals are shown in Figure 5, along with the results obtained under the monotonicity assumption. As all the estimates are very similar, we conclude that the results are robust against violations of the monotonicity assumption.
Interestingly, the estimated direct effects appear to be higher when current smokers are excluded from the data while the indirect effects are largely unchanged. This suggests that the treatment might be less favourable for smokers than the others. The total effect estimates using the simple estimator based on areas under Kaplan-Meier curves resulted in a total effect of $1.9$ years for never smokers or former smokers and negative total effect of $-1.3$ years for current smokers. This comparison, however, is very uncertain as smokers comprised a small group of only 30 individuals.
Figure 5: The lower and upper bounds for the estimates of direct and indirect effects and their $95\%$ credible intervals under relaxing the monotonicity assumption. The estimates in the middle (with identifier ‘M’) are the results obtained under the monotonicity assumption. To the left and right are the lower and upper bounds obtained using the approach of Miles et al. (2017) with either (a) the full sample; (b) dropping smokers from the sample; or (c) using the treatment-dependent confounder model with all one-way interactions between the treatment and each baseline covariate.